OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 24 and the last term is 1,417,176?

Parth (parthkohli):

\[S = a \times {1 - r^n \over 1 - r}\]

Parth (parthkohli):

We already have the value of \(a\); just to find \(r\).

Parth (parthkohli):

There are 5 terms in between term 1 and 6.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {{1,417,176 \over 24} \over 5} }\]

Parth (parthkohli):

That is \(r\). Can you do the rest?

OpenStudy (anonymous):

what is a?

OpenStudy (anonymous):

See you are given last term and first term: a = 24 and \(a_n\) = 1417176 So: \[\large a_n = a \cdot r^{n-1}\] here n = 6 given a = 24 and r = ??? and \(a_n = 1417176\) Can you find r from it??

OpenStudy (anonymous):

i got r = 11,809.8

OpenStudy (anonymous):

Wait...

OpenStudy (anonymous):

\[(24) \cdot r^{6-1} = 1417176\] Can yo tell me what is 6 - 1 = ??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

So what you get in power of r now ??

OpenStudy (anonymous):

\[24 \cdot r^5 = 1417176\] Now divide by 24 both the sides.. And tell me what do you get??

OpenStudy (anonymous):

@emkaye16 I am expecting your reply here.. Simply divide 24 with 24 in LHS.. And divide 1417176 by 24 and tell me what do you get..

OpenStudy (anonymous):

Can you tell me what is : 1417176/24 = ?? You can use calculator no doubt..

OpenStudy (anonymous):

thankyou sorry i did not reply